3 Hundredths + 3 tenths = ______.
 3 Hundredths + 3 tenths = ______.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3 Hundredths + 3 tenths = 0.33.
Place value of the place immediately after the decimal point (i.e. tenth place) is 1/10, that of next place (i.e. hundredths place) is 1/100 and so on.

3 Hundredths is written as = 3 × (1/100)
= 0.03

3 tenths are written as = 3 × (1/10)
= 0.3

Then sum of 3 Hundredths, 3 tenths = 0.03 + 0.3
= 0.33



Thanks